### v2

# The inverse initial data problem for anisotropic Navier–Stokes equations via Legendre time reduction method

Abstract
We consider the inverse initial data problem for the compressible anisotropic Navier-Stokes equations, where the goal is to reconstruct the initial velocity field from lateral boundary observations. This problem arises in applications where direct measurements of internal fluid states are unavailable. We introduce a novel computational framework based on Legendre time reduction, which projects the velocity field onto an exponentially weighted Legendre basis in time. This transformation reduces the original time-dependent inverse problem to a coupled, time-independent elliptic system. The resulting reduced model is solved iteratively using a Picard iteration and a stabilized least-squares formulation under noisy boundary data. Numerical experiments in two dimensions confirm that the method accurately and robustly reconstructs initial velocity fields, even in the presence of significant measurement noise and complex anisotropic structures. This approach offers a flexible and computationally tractable alternative for inverse modeling in fluid dynamics with anisotropic media.
keywords:
inverse initial data problem; anisotropic Navier–Stokes equations; Legendre time reduction
[…]
[label1]organization=Department of Mathematics and Statistics, University of North Carolina at Charlotte,city=Charlotte, postcode=28223, state=NC, country=USA
[…]
[label2]organization=Department of Mathematics, NC State University,city=Raleigh, postcode=27695, state=NC, country=USA
1 Introduction
Let $d\geq 2$ denote the spatial dimension, and let $\Omega\subset\mathbb{R}^{d}$ be a bounded domain with smooth boundary $\partial\Omega$. We consider a fourth-order tensor field $\bm{\mu}:\overline{\Omega}\to\mathbb{R}^{d\times d\times d\times d}$ representing the viscosity property of the medium, which satisfies the coercivity: There exists a constant $\Lambda>0$ such that for all matrices $\bm{\xi}=(\xi_{ij})_{i,j=1}^{d}\in\mathbb{R}^{d\times d}$, we have
$\sum_{i,j,k,l=1}^{d}\mu_{ijkl}\,\xi_{ij}\,\xi_{kl}\geq\Lambda|\bm{\xi}|^{2}.$
Let $T>0$ be a final time. Consider a fluid velocity field $\mathbf{u}:\Omega\times(0,T)\to\mathbb{R}^{d}$ governed by the following initial value problem for the compressible anisotropic Navier–Stokes equations:
$\begin{cases}\displaystyle\frac{\partial\mathbf{u}}{\partial t}+(\mathbf{u}\cdot\nabla)\mathbf{u}=-\nabla p+{\rm div}\left(\bm{\mu}:\nabla\mathbf{u}\right),&({\bf x},t)\in\Omega\times(0,T),\\ {\bf u}({\bf x},t)=0,&({\bf x},t)\in\partial\Omega\times(0,T),\\ \mathbf{u}({\bf x},0)=\mathbf{U}_{0}({\bf x}),&{\bf x}\in\Omega,\end{cases}$ (1.1)
where $\mathbf{U}_{0}:\Omega\to\mathbb{R}^{d}$ represents the initial velocity field and $p:\Omega\times(0,T)\to\mathbb{R}$ is the fluid pressure. Our focus is on the following inverse problem:
Problem 1 (Inverse initial data problem).
Given the lateral boundary observation
$\mathbf{f}({\bf x},t)=\partial_{\nu}\mathbf{u}({\bf x},t),\quad\mbox{for all }({\bf x},t)\in\partial\Omega\times(0,T),$ (1.2)
reconstruct the initial condition $\mathbf{U}_{0}$.
The inverse initial data problem for anisotropic Navier–Stokes equations holds substantial practical importance in a variety of scientific and engineering domains, including seismology, aerospace and automotive engineering, hydraulic systems, oceanography, and atmospheric modeling [24, 34, 37, 38]. In these contexts, direct access to the initial velocity field is often infeasible, while accurate reconstruction of this field is essential for initializing predictive simulations of fluid dynamics. Recovering the initial state from lateral boundary measurements thus constitutes a critical and timely inverse problem. From a theoretical standpoint, this problem presents significant challenges. The well-posedness of the forward anisotropic Navier–Stokes system remains unresolved in general settings, particularly in three dimensions and without strong regularity assumptions. For the inverse problem, the absence of Carleman estimates or other analytical tools applicable to fully anisotropic fourth-order tensors makes it difficult to establish uniqueness or stability. As a result, many foundational theoretical questions remain open in this setting.
In this work, we focus exclusively on the computational aspect of the problem. We do not attempt to address the underlying analytical difficulties but instead develop a numerical framework based on time-dimensional reduction and iterative regularization. The proposed method achieves stable and accurate reconstructions in practice, providing a viable computational approach for a class of inverse problems that currently lack rigorous theoretical guarantees.
While the primary focus of this paper is on an inverse problem associated with the anisotropic Navier–Stokes equations, it is essential first to understand the status of the corresponding forward problem. The well-posedness and regularity of the forward model form the foundation upon which the theoretical analysis and numerical treatment of inverse problems are built. Therefore, we briefly review relevant results on the forward problem. While the global existence and regularity of solutions to the Navier–Stokes equations remain one of the most challenging open problems in mathematical fluid dynamics, substantial progress has been made in specific settings. Foundational analytical frameworks have been developed for various formulations of the Navier–Stokes system, particularly in the compressible and stationary cases [4, 16, 24, 29, 35]. The existence of global weak solutions for stationary problems with homogeneous Dirichlet boundary conditions was first established in [24], and later refined and generalized in [29, 6, 32, 33]. Further developments in the well-posedness theory include the treatment of compressible stationary Navier–Stokes equations with small data, where uniqueness and differentiability were proven in [31]. More recent advances have addressed coupled systems and complex boundary interactions: global well-posedness for a Navier–Stokes–Cahn–Hilliard model with chemotaxis and singular potential in two dimensions was established in [8], while local and global results for the Boussinesq system with viscoelastic effects under various boundary conditions were obtained in [2]. In three dimensions, global existence and decay properties of strong solutions to the nonhomogeneous incompressible Navier–Stokes–Cahn–Hilliard system were demonstrated in [3] using a priori energy methods. Finally, the well-posedness of stationary Navier–Stokes equations in higher-dimensional half-spaces was studied in [7], where scaling-critical Besov spaces and maximal regularity theory were employed to derive existence and asymptotic behavior results.
Inverse problems for the Navier–Stokes equations have been extensively studied to determine unknown coefficients or initial data from boundary or interior measurements. In [5], a numerical approximation for inverse problems was developed using a Lagrange–Galerkin method that effectively addresses drag reduction and velocity identification. A regularized Gauss–Newton method for shape reconstruction in two-dimensional Navier–Stokes equations was derived in [39], establishing differentiability with respect to the boundary curve. Global identifiability of viscosity in two-dimensional Stokes and Navier–Stokes equations using Cauchy force measurements was proved in [17], advancing prior Dirichlet-to-Neumann map approaches. Global uniqueness for inverse boundary value problems for the Navier–Stokes equations and Lamé system in two dimensions was established in [10] using complex geometric optics solutions. Iterative methods, including a novel simple backward integration approach, were introduced in [30] to solve retrospective inverse problems, outperforming direct adjoint looping. Global uniqueness for viscosity determination in Stokes and Navier–Stokes equations in two and three dimensions was demonstrated in [25], combining prior theorems with new results for arbitrary bounded domains.
Our strategy for solving the inverse problem associated with the Navier–Stokes equations stated in Problem 1 is grounded in the use of the Legendre time-reduction technique. Specifically, we approximate the velocity field $\mathbf{u}$ in the time variable by truncating its Fourier series with respect to an exponentially weighted Legendre basis of the weighted space $L_{e^{-2t}}^{2}(0,T)$. This basis, recently introduced in [36] for inverse initial data problems in elasticity, and further utilized in [23] for Maxwell’s system, offers notable advantages over classical bases such as the trigonometric system and the Legendre polynomial without the exponential weight. In particular, the exponential-weighted Legendre polynomials possess two desirable properties: non-vanishing derivatives and favorable convergence behavior of the truncated series applied to time derivatives of the velocity field. By applying this time-reduction procedure, we reformulate the original time-dependent inverse problem into a sequence of time-independent systems, where the principal operator takes the form ${\rm div}(\bm{\mu}:\nabla\cdot)$. We refer to this reformulated system as the time-dimensional reduction model. Leveraging this structure, we then employ a Picard iteration to solve this model. The obtained solution directly yields an approximate solution to the original inverse problem, thereby enabling a globally convergent reconstruction scheme.
The remainder of the paper is organized as follows. Section 2 reviews the exponentially weighted Legendre basis used for temporal approximation. Section 3 derives the time-dimensional reduction model that converts the time-dependent inverse problem into a coupled elliptic system. Section 4 presents a Picard, least-squares (quasi-reversibility) algorithm for solving the reduced model. Section 5 reports numerical experiments demonstrating accuracy and robustness. We then include a dedicated Section 6 on the stability of the time reduction and the choice of time basis, including a scaled weight for large final times $T$. Section 7 concludes with a discussion and future directions.
2 The Legendre polynomial-exponential basis
The Legendre polynomial-exponential basis, first introduced in [36], is fundamental to the time-dimensional reduction method employed in this study. This basis integrates the spectral characteristics of classical Legendre polynomials with an exponential weight, providing a structured approximation framework in the space of exponentially weighted square-integrable functions. For clarity and completeness, we briefly revisit this basis below. Let $\{P_{n}\}_{n\geq 0}$ denote the Legendre polynomials on (1,1)(-1,1), defined by Rodrigues’ formula:
$P_{n}(x)=\frac{1}{2^{n}n!}\frac{d^{n}}{dx^{n}}(x^{2}-1)^{n}.$
To define an orthonormal basis on $(0,T)$, we introduce the affine transformation $x=\frac{2t}{T}-1$, and define the rescaled polynomials
$Q_{n}(t):=\sqrt{\frac{2n+1}{T}}P_{n}\left(\frac{2t}{T}-1\right)\quad\mbox{for }t\in(0,T).$
The set $\{Q_{n}\}_{n\geq 0}$ forms an orthonormal basis in $L^{2}(0,T)$. We then define the Legendre polynomial-exponential basis by:
$\Psi_{n}(t):=e^{t}Q_{n}(t)\quad\mbox{for }t\in(0,T).$
The system $\{\Psi_{n}\}_{n\geq 0}$ forms an orthonormal basis of the exponentially weighted space
$L^{2}_{e^{-2t}}(0,T):=\left\{u\in L^{2}(0,T)\,\middle|\,\int_{0}^{T}e^{-2t}|u(t)|^{2}dt<\infty\right\},$
with the inner product
$\langle u,v\rangle_{L^{2}_{e^{-2t}}(0,T)}:=\int_{0}^{T}e^{-2t}u(t)v(t)\,dt.$
Proposition 1.
The Legendre polynomial-exponential basis functions $\Psi_{n}$, for $n\geq 0$, satisfy the following regularity and spectral properties:
  1. 1.
    (Smoothness and non-vanishing derivatives) For each $n\geq 0$, the function $\Psi_{n}$ is infinitely differentiable on $(0,T)$, and none of its derivatives of any order vanishes identically on this interval.
  2. 2.
    (Spectral coefficient decay) For every integer $k\in\mathbb{N}$, there exists a constant $C>0$, depending only on $k$ and $T$, such that for all $u\in H^{k}(0,T)$,
    $\sum_{n=0}^{\infty}n^{2k}\left|\langle u,\Psi_{n}\rangle_{L^{2}_{e^{-2t}}(0,T)}\right|^{2}\leq C\|u\|_{H^{k}(0,T)}^{2}.$ (2.1)
  3. 3.
    (Derivative norm growth) There exists a constant $C>0$, depending only on $T$, such that for all $n\geq 1$,
    $\|\Psi_{n}^{\prime}\|_{L^{2}_{e^{-2t}}(0,T)}\leq Cn^{3/2},\quad\text{and}\quad\|\Psi_{n}^{\prime\prime}\|_{L^{2}_{e^{-2t}}(0,T)}\leq Cn^{7/2}.$ (2.2)
Remark 1 (The importance of the weight $e^{t}$).
We omit the proof of Proposition 1, which follows from Proposition 2.1, Lemma 2.1, and Lemma 2.2 in [36]. Note that Item 1 relies on the factor $e^{t}$ in $\Psi_{n}=e^{t}Q_{n}$: without this factor, derivatives of some $\Psi_{n}$ may be identically zero (e.g., the constant time mode), which can reduce the effectiveness of the time-dimensional reduction; see Subsection 6.2 for details.
Let $N$ be a positive integer. We introduce three key operators associated with this basis.
  • Fourier Coefficient Operator ($\mathbb{F}^{N}$): For $u\in L^{2}_{e^{-2t}}\left((0,T);L^{2}(\Omega)^{d}\right)$, we define the vector of its first $N+1$ coefficients by:
    $\mathbb{F}^{N}[u]({\bf x}):=\begin{bmatrix}u_{0}({\bf x})&u_{1}({\bf x})&\cdots&u_{N}({\bf x})\end{bmatrix}^{\top}$,
    where
    $u_{n}({\bf x}):=\int_{0}^{T}e^{-2t}u({\bf x},t)\Psi_{n}(t)dt$
    for ${\bf x}\in\Omega.$
  • Expansion Operator ($\mathbb{S}^{N}$): Given the coefficient vector
    $\mathbf{U}=(u_{0},\ldots,u_{N})^{\top}\in L^{2}(\Omega)^{d\times(N+1)},$
    the space-time expansion is defined by
    $\mathbb{S}^{N}[\mathbf{U}]({\bf x},t):=\sum_{n=0}^{N}u_{n}({\bf x})\Psi_{n}(t),$
    for $({\bf x},t)\in\Omega\times(0,T).$
  • Projection Operator ($\mathbb{P}^{N}$): The projection of $u\in L^{2}_{e^{-2t}}((0,T);L^{2}(\Omega)^{d})$ onto the subspace $\mathbb{V}^{N}:=\mathrm{span}\{\Psi_{0},\ldots,\Psi_{N}\}$ is:
    $\mathbb{P}^{N}[u]({\bf x},t):=\sum_{n=0}^{N}\left(\int_{0}^{T}e^{-2t}u({\bf x},t)\Psi_{n}(t)dt\right)\Psi_{n}(t),$ (2.3)
    for $({\bf x},t)\in\Omega\times(0,T).$
These operators satisfy the identity
$\mathbb{S}^{N}[\mathbb{F}^{N}[u]]=\mathbb{P}^{N}[u],$
and Parseval’s identity implies:
$\|\mathbb{P}^{N}[u]\|_{L^{2}_{e^{-2t}}((0,T);L^{2}(\Omega)^{d})}=\|\mathbf{U}\|_{L^{2}(\Omega)^{d\times(N+1)}}\leq\|u\|_{L^{2}_{e^{-2t}}((0,T);L^{2}(\Omega)^{d})}.$ (2.4)
Remark 2.
For any integer $N\geq 1$, the operators $\mathbb{F}^{N}$, $\mathbb{S}^{N}$, and $\mathbb{P}^{N}$ admit natural restrictions to vector-valued functions in the space $L^{2}_{e^{-2t}}\left((0,T);H^{p}(\Omega)^{d}\right)$ for any $p\geq 1$. In this setting, the Parseval-type inequality in (2.4) continues to hold with $L^{2}(\Omega)^{d}$ replaced by $H^{p}(\Omega)^{d}$. Explicitly, we have
$\|\mathbb{P}^{N}[u]\|_{L^{2}_{e^{-2t}}((0,T);H^{p}(\Omega)^{d})}=\|\mathbf{U}\|_{H^{p}(\Omega)^{d\times(N+1)}}\leq\|u\|_{L^{2}_{e^{-2t}}((0,T);H^{p}(\Omega)^{d})}.$
3 The time-dimensional reduction method
The inverse problem stated in Problem 1 can be reformulated as the recovery of a vector field $\mathbf{u}$ satisfying
$\begin{cases}\partial_{t}\mathbf{u}+(\mathbf{u}\cdot\nabla)\mathbf{u}=-\nabla p+{\rm div}(\bm{\mu}:\nabla\mathbf{u})&\text{in }\Omega\times(0,T),\\ \mathbf{u}({\bf x},t)=\textbf{0}&\text{on }\partial\Omega\times(0,T),\\ \partial_{\nu}\mathbf{u}({\bf x},t)=\mathbf{f}({\bf x},t)&\text{on }\partial\Omega\times(0,T),\end{cases}$ (3.1)
given the data $\mathbf{f}$ as in (1.2).
Remark 3.
The existence of a solution to (3.1) is guaranteed, as $\mathbf{u}$ corresponds to the restriction of the solution to the forward problem (1.1) on $\Omega\times(0,T)$. Although a rigorous proof of uniqueness for the original inverse problem is beyond the scope of this work, which is primarily concerned with computational methodologies, we emphasize that this question remains of significant theoretical interest.
We now derive the time-dimensional reduction model. Let $N\geq 1$ be a fixed integer, and consider the projection operator $\mathbb{P}^{N}$. Applying $\mathbb{P}^{N}$ to both sides of the differential equation in (3.1), we obtain:
$\mathbb{P}^{N}[\partial_{t}\mathbf{u}^{*}]+\mathbb{P}^{N}[(\mathbf{u}\cdot\nabla)\mathbf{u}]=-\mathbb{P}^{N}[\nabla p]+{\rm div}(\bm{\mu}:\nabla\mathbb{P}^{N}[\mathbf{u}])\quad\mbox{in }\Omega\times(0,T).$ (3.2)
Adding and subtracting $\partial_{t}\mathbb{P}^{N}[{\bf u}]+(\mathbb{P}^{N}[{\bf u}]\cdot\nabla)\mathbb{P}^{N}[{\bf u}]$ on the left-hand side of (3.2) and then rearranging the resulting equation yields
$\partial_{t}\mathbb{P}^{N}[\mathbf{u}]-{\rm div}(\bm{\mu}:\nabla\mathbb{P}^{N}[\mathbf{u}])+(\mathbb{P}^{N}[\mathbf{u}]\cdot\nabla)\mathbb{P}^{N}[\mathbf{u}]+\mathbb{P}^{N}[\nabla p]\\ =-\left(\mathbb{P}^{N}[\partial_{t}\mathbf{u}]-\partial_{t}\mathbb{P}^{N}[\mathbf{u}]\right)-\left(\mathbb{P}^{N}[(\mathbf{u}\cdot\nabla)\mathbf{u}]-(\mathbb{P}^{N}[\mathbf{u}]\cdot\nabla)\mathbb{P}^{N}[\mathbf{u}]\right)$ (3.3)
in $\Omega\times(0,T)$. We have the theorem
Theorem 1 (Asymptotic commutation theorem).
Assume the solution $\bf u$} to (3.1) belongs to the class $H^{3}((0,T),H^{2}(\Omega)^{d})$. Then,
$\lim_{N\to\infty}\left\|\mathbb{P}^{N}[\partial_{t}\mathbf{u}]-\partial_{t}\mathbb{P}^{N}[\mathbf{u}]\right\|_{L^{2}((0,T);H^{2}(\Omega)^{d})}=0$ (3.4)
and
$\lim_{N\to\infty}\left\|\mathbb{P}^{N}[(\mathbf{u}\cdot\nabla)\mathbf{u}]-(\mathbb{P}^{N}[\mathbf{u}]\cdot\nabla)\mathbb{P}^{N}[\mathbf{u}]\right\|_{H^{2}(\Omega)^{d}}=0.$ (3.5)
Proof.
We first establish the convergence of the series $\sum_{n=0}^{\infty}\left\langle{\bf u}({\bf x},\cdot),\Psi_{n}\right\rangle_{L^{2}_{e^{-2t}}(0,T)}\Psi_{n}^{\prime}(t)$ in $L^{2}(\Omega)^{d}.$ It follows from (2.1) for the vector $\bf u$} and for $k=3$ that for each ${\bf x}\in\Omega$
$\sum_{n=0}^{\infty}n^{6}\left|\langle{\bf u}({\bf x},\cdot),\Psi_{n}\rangle_{L^{2}_{e^{-2t}}(0,T)^{d}}\right|^{2}\leq C\|{\bf u}({\bf x},\cdot)\|_{H^{3}(0,T)^{d}}^{2}.$ (3.6)
Since each term of the series in the left-hand side of (3.6) is non-negative, we have for all $n\geq 1$,
$n^{6}\left|\langle{\bf u}({\bf x},\cdot),\Psi_{n}\rangle_{L^{2}_{e^{-2t}}(0,T)^{d}}\right|^{2}\leq C\|{\bf u}({\bf x},\cdot)\|_{H^{3}(0,T)^{d}}^{2}.$ (3.7)
Integrating both sides of (3.7) gives
$n^{6}\left\|\langle{\bf u}(\cdot,\cdot),\Psi_{n}\rangle_{L^{2}_{e^{-2t}}(0,T)}\right\|_{L^{2}(\Omega)^{d}}^{2}\leq C\int_{\Omega}\|{\bf u}({\bf x},\cdot)\|_{H^{3}(0,T)}^{2}d{\bf x}.$ (3.8)
Therefore, multiplying $e^{-2t}|\Psi_{n}^{\prime}(t)|^{2}$ to both sides of (3.8) and integrating the resulting inequality, we obtain
$\int_{0}^{T}n^{6}e^{-2t}\left\|\langle{\bf u}(\cdot,\cdot),\Psi_{n}\rangle_{L^{2}_{e^{-2t}}(0,T)}\right\|_{L^{2}(\Omega)^{d}}^{2}|\Psi^{\prime}(t)|^{2}dt$ $\leq C\int_{0}^{T}e^{-2t}\int_{\Omega}\|{\bf u}({\bf x},\cdot)\|_{H^{3}(0,T)}^{2}|\Psi^{\prime}(t)|^{2}d{\bf x}dt$
$\leq C\int_{\Omega}\|{\bf u}({\bf x},\cdot)\|_{H^{3}(0,T)}^{2}d{\bf x}\int_{0}^{T}e^{-2t}|\Psi^{\prime}_{n}(t)|^{2}dt$
$\leq Cn^{3}\int_{\Omega}\|{\bf u}({\bf x},\cdot)\|_{H^{3}(0,T)}^{2}d{\bf x}.$
Here, we have used (2.2) at the last estimate. Hence,
$\int_{0}^{T}e^{-2t}\left\|\langle{\bf u}(\cdot,\cdot),\Psi_{n}\rangle_{L^{2}_{e^{-2t}}(0,T)}\right\|_{L^{2}(\Omega)^{d}}^{2}|\Psi^{\prime}(t)|^{2}dt\leq Cn^{-3}\int_{\Omega}\|{\bf u}({\bf x},\cdot)\|_{H^{3}(0,T)}^{2}d{\bf x}$ (3.9)
for $n\geq 1$. Here, we avoid the case $n=0$ because of the appearance of $n^{-3}$ on the right-hand side of (3.9). It follows from (3.9) that
$\sum_{n=1}^{\infty}\Big(\int_{0}^{T}e^{-2t}\left\|\langle{\bf u}(\cdot,\cdot),\Psi_{n}\rangle_{L^{2}_{e^{-2t}}(0,T)}\right\|_{L^{2}(\Omega)^{d}}^{2}|\Psi^{\prime}(t)|^{2}dt\Big)^{1/2}\\ \leq C\sum_{n=1}^{\infty}n^{-3/2}\Big(\int_{\Omega}\|{\bf u}({\bf x},\cdot)\|_{H^{3}(0,T)}^{2}d{\bf x}\Big)^{1/2}<\infty.$ (3.10)
It follows from (3.10) that the series $\sum_{n=0}^{\infty}\langle{\bf u}({\bf x},\cdot),\Psi_{n}\rangle_{L^{2}_{e^{-2t}}(0,T)}\Psi_{n}^{\prime}(t)$ Define the limit function
${\bf w}({\bf x},t):=\lim_{N\to\infty}\sum_{n=0}^{N}\langle{\bf u}(\cdot,\cdot),\Psi_{n}\rangle_{L^{2}_{e^{-2t}}(0,T)}\,\Psi_{n}^{\prime}(t).$
We show that ${\bf w}={\bf u}_{t}$. Let $\varphi\in C_{c}^{\infty}(\Omega\times(0,T))^{d}$ be an arbitrary test function. We compute:
$\int_{\Omega}\int_{0}^{T}{\bf w}({\bf x},t)\cdot\varphi({\bf x},t)\,{\rm d}{\bf x}{\rm d}t$ $=\int_{\Omega}\int_{0}^{T}\lim_{N\to\infty}\left(\sum_{n=0}^{N}\langle{\bf u}({\bf x},\cdot),\Psi_{n}\rangle_{L^{2}_{e^{-2t}}(0,T)}\Psi_{n}^{\prime}(t)\right)\cdot\varphi({\bf x},t)\,{\rm d}{\bf x}{\rm d}t$
$=\lim_{N\to\infty}\sum_{n=0}^{N}\int_{\Omega}\int_{0}^{T}\langle{\bf u}({\bf x},\cdot),\Psi_{n}\rangle_{L^{2}_{e^{-2t}}(0,T)}\,\Psi_{n}^{\prime}(t)\cdot\varphi({\bf x},t)\,{\rm d}{\bf x}{\rm d}t.$
This interchange of the limit and the integral above is justified by the convergence with respect to the norm of $L^{2}_{e^{-2t}}((0,T);L^{2}(\Omega)^{d})$ of the partial sums to $\bf w$}, while $\varphi$ is compactly supported and bounded. Next, we integrate by parts in time:
$\lim_{N\to\infty}\sum_{n=0}^{N}\int_{\Omega}\int_{0}^{T}\langle{\bf u}({\bf x},\cdot),\Psi_{n}$ $\rangle_{L^{2}_{e^{-2t}}(0,T)}\Psi_{n}^{\prime}(t)\cdot\varphi({\bf x},t)\,{\rm d}{\bf x}{\rm d}t$
$=\lim_{N\to\infty}\sum_{n=0}^{N}\int_{\Omega}\langle{\bf u}({\bf x},\cdot),\Psi_{n}\rangle_{L^{2}_{e^{-2t}}(0,T)}\cdot\left(\int_{0}^{T}\Psi_{n}^{\prime}(t)\varphi({\bf x},t){\rm d}t\right){\rm d}{\bf x}$
$=-\int_{\Omega}\lim_{N\to\infty}\sum_{n=0}^{N}\langle{\bf u}({\bf x},\cdot),\Psi_{n}\rangle_{L^{2}_{e^{-2t}}(0,T)}\cdot\left(\int_{0}^{T}\Psi_{n}(t)\varphi_{t}({\bf x},t){\rm d}t\right){\rm d}{\bf x}$
$=-\int_{\Omega}\int_{0}^{T}{\bf u}({\bf x},t)\cdot\varphi_{t}({\bf x},t)\,{\rm d}{\bf x}{\rm d}t,$
where we used completeness of the $\{\Psi_{n}\}$ basis in $L^{2}_{e^{-2t}}(0,T)$, convergence of the series, and the dominated convergence theorem again. Hence, ${\bf w}={\bf u}_{t}$ in the distributional sense, and since both of these vector-valued functions lie in $L^{2}_{e^{-2t}}((0,T);L^{2}(\Omega)^{d})$, the equality holds almost everywhere.
We can repeat this arguments for $\partial_{x_{i}}{\bf u}({\bf x},t)$ and $\partial_{x_{i}x_{j}}{\bf u}({\bf x},t)$, $1\leq i,j\leq d$ to obtain the convergences of $\sum_{n=0}^{\infty}\langle\partial_{x_{i}}{\bf u}({\bf x},\cdot),\Psi_{n}\rangle_{L^{2}_{e^{-2t}}(0,T)}\Psi_{n}^{\prime}(t)$ and $\sum_{n=0}^{\infty}\langle\partial_{x_{i}x_{j}}{\bf u}({\bf x},\cdot),\Psi_{n}\rangle_{L^{2}_{e^{-2t}}(0,T)}\Psi_{n}^{\prime}(t)$ to $\partial_{x_{i}}{\bf w}$ and $\partial_{x_{i}x_{j}}{\bf w}$ respectively in $L^{2}_{e^{-2t}}((0,T);L^{2}(\Omega)^{d})$. We obtain (3.4).
Proving (3.5) is straightforward because the projection operator $\mathbb{P}^{N}$ acts only on the temporal variable $t$ and does not affect the spatial variable $\bf x$}. ∎
Remark 4 (Regularity in Theorem 3.5 and observed performance).
Theorem 3.5 assumes high regularity (for example, $\mathbf{u}\in H^{3}(0,T;H^{2}(\Omega)^{d})$). These assumptions are used to carry out the analysis and obtain clear statements with explicit bounds.
In computations, the algorithm performs well even when the initial condition $\mathbf{u}_{0}$ is only in $L^{2}(\Omega)^{d}$ and may be discontinuous, so the hypotheses of Theorem 3.5 do not hold, see Section 5 for our out-of-expectation numerical results. This indicates that the method’s practical scope is broader than current theory covers.
Due to Theorem 3.5 and the limits (3.4) and (3.5), equation (3.3) can be approximated by the reduced system
$\partial_{t}\mathbb{P}^{N}[\mathbf{u}]+(\mathbb{P}^{N}[\mathbf{u}]\cdot\nabla)\mathbb{P}^{N}[\mathbf{u}]=-\mathbb{P}^{N}[\nabla p]+{\rm div}(\bm{\mu}:\nabla\mathbb{P}^{N}[\mathbf{u}]),$ (3.11)
which serves as the foundation of our time-dimensional reduction model. For each $n\geq 0$, let
$\mathbf{u}_{n}({\bf x})=\int_{0}^{T}e^{-2t}\mathbf{u}({\bf x},t)\Psi_{n}(t)\,dt,$ (3.12)
be the $n$-th Fourier mode of $\mathbf{u}$ with respect to the basis $\{\Psi_{n}\}_{n\geq 0}$.
Remark 5 (Asymptotic commutation).
For large $N$, due to Theorem 3.5, the spectral projection $\mathbb{P}^{N}$ approximatedly commutes with $\partial_{t}$ and with $({\bf u}\cdot\nabla){\bf u}$. In other words, “project-then-operate” and “operate-then-project” agree in the limit. This consideration motivates our choice of the exponentially weighted Legendre basis $\{\Psi_{n}\}_{n\geq 0}$. Other time bases (e.g., the exponentially weighted polynomial basis in $L^{2}(0,T)$; see [13, 28]) need not have this property and may therefore be unsuitable for deriving (3.12).
Due to (2.3),
$\mathbb{P}^{N}[\mathbf{u}]({\bf x},t)=\sum_{n=0}^{N}\mathbf{u}_{n}({\bf x})\Psi_{n}(t).$ (3.13)
Substituting (3.13) and into the system (3.11), we obtain the following approximation:
$\sum_{n=0}^{N}\mathbf{u}_{n}({\bf x})\Psi_{n}^{\prime}(t)+\sum_{l=0}^{N}\sum_{n=0}^{N}\left(\mathbf{u}_{l}({\bf x})\cdot\nabla\right)\mathbf{u}_{n}({\bf x})\Psi_{l}(t)\Psi_{n}(t)=-\nabla p+\sum_{n=0}^{N}{\rm div}\left(\bm{\mu}:\nabla\mathbf{u}_{n}({\bf x})\right)\Psi_{n}(t),$ (3.14)
valid for all $({\bf x},t)\in\Omega\times(0,T)$. To derive the time-reduced model, we test equation (3.14) against $e^{-2t}\Psi_{m}(t)$ and integrate over $t\in(0,T)$ for each $m=0,1,\dots,N$, resulting in:
$\sum_{n=0}^{N}s_{mn}\mathbf{u}_{n}({\bf x})+\sum_{l=0}^{N}\sum_{n=0}^{N}a_{mnl}\left(\mathbf{u}_{l}({\bf x})\cdot\nabla\right)\mathbf{u}_{n}({\bf x})=-\nabla p_{m}({\bf x})+{\rm div}\left(\bm{\mu}:\nabla\mathbf{u}_{m}({\bf x})\right),$ (3.15)
for all ${\bf x}\in\Omega$, where the coupling coefficients are defined by
$s_{mn}=\int_{0}^{T}e^{-2t}\Psi_{n}^{\prime}(t)\Psi_{m}(t)\,dt,\quad a_{mnl}=\int_{0}^{T}e^{-2t}\Psi_{l}(t)\Psi_{n}(t)\Psi_{m}(t)\,dt,$
and
$p_{m}({\bf x})=\int_{0}^{T}e^{-2t}p({\bf x},t)\Psi_{m}(t)dt.$
The corresponding Cauchy boundary conditions for each mode $\mathbf{u}_{m}$ follow from (1.2) and (3.12):
$\mathbf{u}_{m}({\bf x})=\textbf{0}\quad\partial_{\nu}\mathbf{u}_{m}({\bf x})=\int_{0}^{T}e^{-2t}\mathbf{f}({\bf x},t)\Psi_{m}(t)\,dt,$ (3.16)
for all ${\bf x}\in\partial\Omega$. Equations (3.15) and (3.16) define a coupled boundary value problem for the vector of coefficients $\mathbb{F}^{N}[\mathbf{u}]$. We refer to this as the time-dimensional reduction model:
$\begin{cases}\displaystyle\sum_{n=0}^{N}s_{mn}\mathbf{u}_{n}({\bf x})+\sum_{l=0}^{N}\sum_{n=0}^{N}a_{mnl}\left(\mathbf{u}_{l}({\bf x})\cdot\nabla\right)\mathbf{u}_{n}({\bf x})=-\nabla p_{m}+{\rm div}\left(\bm{\mu}:\nabla\mathbf{u}_{m}({\bf x})\right),&{\bf x}\in\Omega,\\ \displaystyle\mathbf{u}_{m}({\bf x})=\textbf{0},&{\bf x}\in\partial\Omega,\\ \displaystyle\partial_{\nu}\mathbf{u}_{m}({\bf x})=\mathbf{f}_{m}({\bf x}):=\int_{0}^{T}e^{-2t}\mathbf{f}({\bf x},t)\Psi_{m}(t)\,dt,&{\bf x}\in\partial\Omega,\end{cases}$ (3.17)
for $m=0,1,\dots,N$.
Solving the system (3.17) yields the vector of Fourier modes $\mathbb{F}^{N}[\mathbf{u}]$, which can be used to reconstruct an approximation of the original space-time solution $\mathbf{u}({\bf x},t)$ via the expansion $\mathbb{P}^{N}[\mathbf{u}]=\mathbb{S}^{N}[\mathbb{F}^{N}[\mathbf{u}]]$. According to the approximation result in (3.13), this reconstruction provides a full space-time representation of $\mathbf{u}$, thereby transforming the inverse problem under consideration (Problem 1) into the computational task of solving the reduced model (3.17). Despite this reduction, solving the system (3.17) remains highly nontrivial due to the presence of the nonlinear convective term
$\sum_{l=0}^{N}\sum_{n=0}^{N}a_{mnl}\left(\mathbf{u}_{l}({\bf x})\cdot\nabla\right)\mathbf{u}_{n}({\bf x}),\quad m\in\{0,1,\dots,N\},$ (3.18)
making (3.17) a coupled nonlinear system. We will apply a conventional Picard iteration to overcome this difficulty, see Section 4.
Remark 6.
The time-dimensional reduction model (3.17) is an approximate representation of the original system, stemming from the truncation of the infinite Fourier expansion in (3.13). Despite this, we adopt the working assumption that
$\mathbf{U}=\begin{bmatrix}\mathbf{u}_{0}&\mathbf{u}_{1}&\dots&\mathbf{u}_{N}\end{bmatrix}^{\top}$
satisfies the reduced model. This assumption is justified by the fact that the truncation error diminishes as the cutoff parameter $N$ increases, with the error terms in the left-hand sides of (3.4) and (3.5) vanishing in the limit as $N\to\infty$, see Theorem 3.5. Consequently, the reduced model becomes an increasingly accurate surrogate for the full dynamics. Moreover, from a computational standpoint, the assumption is natural: practical implementations inherently involve a finite $N$, making the exact infinite-dimensional formulation intractable.
Remark 7.
The uniqueness of the time-reduction model (3.17) remains an open problem due to the complexity introduced by the anisotropic fourth-order tensor $\bm{\mu}$ and the presence of the nonlinear convection term given in (3.18). A well-established approach in the literature for establishing uniqueness in inverse problems is the use of Carleman estimates. However, deriving a Carleman estimate in the fully anisotropic setting is highly nontrivial and currently lacks a general theoretical framework. In contrast, if the problem is restricted to the isotropic case, Carleman estimates become more tractable; see, for instance, [9, 11]. Under such simplifications, the uniqueness of the time-reduction model (3.17) could then be rigorously analyzed. Therefore, we reserve a detailed investigation of uniqueness and stability results based on Carleman estimates for the isotropic case as a direction for future research.
4 A Picard algorithm to solve the time-reduction model
We define the admissible set
$H:=\left\{\mathbf{U}\in H^{2}(\Omega)^{d\times(N+1)}\ \middle|\ \mathbf{U}|_{\partial\Omega}=\mathbf{0}\right\},$
which contains the solution of the reduced system in (3.17). Let the initial guess of the true solution be denoted by
${\bf U}^{(0)}=\begin{bmatrix}{\bf u}^{(0)}_{0}&{\bf u}^{(0)}_{1}&\dots&{\bf u}^{(0)}_{N}\end{bmatrix}^{\top}$.
Assuming
${\bf U}^{(k)}=\begin{bmatrix}{\bf u}^{(k)}_{0}&{\bf u}^{(k)}_{1}&\dots&{\bf u}^{(k)}_{N}\end{bmatrix}^{\top}$
is known for some $k\geq 0$, we define the next iterate
${\bf U}^{(k+1)}=\begin{bmatrix}{\bf u}^{(k+1)}_{0}&{\bf u}^{(k+1)}_{1}&\dots&{\bf u}^{(k+1)}_{N}\end{bmatrix}^{\top}$
as the solution to the following boundary value problem:
$\begin{cases}\displaystyle\sum_{n=0}^{N}s_{mn}\mathbf{u}_{n}^{(k+1)}({\bf x})+\sum_{l=0}^{N}\sum_{n=0}^{N}a_{mnl}\left(\mathbf{u}_{l}^{(k)}({\bf x})\cdot\nabla\right)\mathbf{u}_{n}^{(k)}({\bf x})\\ \hskip 142.26378pt=-\nabla p_{m}+{\rm div}\left(\bm{\mu}:\nabla\mathbf{u}_{m}^{(k+1)}({\bf x})\right),&{\bf x}\in\Omega,\\ \mathbf{u}_{m}^{(k+1)}({\bf x})=\mathbf{0},&{\bf x}\in\partial\Omega,\\ \partial_{\nu}\mathbf{u}_{m}^{(k+1)}({\bf x})=\mathbf{f}_{m}({\bf x}),&{\bf x}\in\partial\Omega,\end{cases}$ (4.1)
for each $m=0,1,\dots,N$. This iteration scheme is inspired by the classical Picard method. In general, the convergence of the sequence $\{{\bf U}^{(k)}\}_{k\geq 0}$ depends on the proximity of the initial guess ${\bf U}^{(0)}$ to the true solution of (3.17). However, we do not pursue this local convergence analysis, as such accurate initial guesses are typically unavailable in practice. Instead, we demonstrate numerically that the iteration converges even when the initial guess is chosen as ${\bf U}^{(0)}=\mathbf{0}$.
The next step is to solve the over-determined problem (4.1). By over-determined, we mean the presence of both Dirichlet and Neumann information of ${\bf U}^{(k+1)}$ in this problem. We propose to use the quasi-reversibility method, first introduced in [18]. That means the solution ${\bf U}^{(k+1)}$ is the minimizer of a least-squares functional involving a Tikhonov regularization term
$J^{\epsilon,N}_{{\bf U}^{(k)}}(\bm{\varphi})=\sum_{m=0}^{N}\Bigg[\int_{\Omega}\Big|{\rm div}\left(\bm{\mu}:\nabla\varphi_{m}({\bf x})\right)-\sum_{n=0}^{N}s_{mn}\varphi_{n}({\bf x})-\sum_{l=0}^{N}\sum_{n=0}^{N}a_{mnl}\left({\bf u}_{l}^{(k)}({\bf x})\cdot\nabla\right)\mathbf{v}_{n}^{(k)}({\bf x})\\ -\nabla p_{m}({\bf x})\Big|^{2}d{\bf x}+\int_{\partial\Omega}|\partial_{\nu}\varphi_{m}-\mathbf{f}_{m}|^{2}d\sigma({\bf x})+\epsilon\|\varphi_{m}\|_{H^{2}(\Omega)^{d}}^{2}\Bigg].$ (4.2)
The existence and uniqueness of the minimizer of $J^{\epsilon,N}_{{\bf U}^{(k)}}$ are guaranteed by the strict convexity of the functional and the fact that its domain $H$ is a Hilbert space.
We summarize the Picard iteration and quasi-reversibility method to solve (3.17) and hence Problem 1 in Algorithm 1.
Algorithm 1 Solution procedure for Problem 1 via the time dimension reduction method
1:Select a cutoff number NN the regularization parameter $\epsilon$.
2:Set ${\bf U}^{(0)}={\bf 0}\in H.$
3:for $k=0$ to K1K-1, for some $K\in\mathbb{N}$ do
4:  Minimize $J^{\epsilon,N}_{{\bf U}^{(k)}}$ defined in (4.2) and set ${\bf U}^{(k+1)}$ as the obtained minimizer.
5:end for
6:Set the computed solution ${\bf U}^{\rm comp}={\bf U}^{K}$:
$U^{\rm comp}=\begin{bmatrix}{\bf u}_{0}^{\rm comp}&{\bf u}_{1}^{\rm comp}&\dots&{\bf u}_{N}^{\rm comp}\end{bmatrix}^{\top}$.
7:Reconstruct the fluid velocity field ${\bf u}^{\rm comp}$ using space-time expansion:
${\bf u}^{\rm comp}(\mathbf{x},t)=\mathbb{S}^{N}[U^{\rm comp}]=\sum_{n=0}^{N}{\bf u}_{n}^{\rm comp}(\mathbf{x})\Psi_{n}(t),$ (4.3)
for all $(\mathbf{x},t)\in\Omega\times(0,T)$.
8:Compute the approximate initial velocity field:
$\mathbf{U}_{0}^{\rm comp}(\mathbf{x})={\bf u}^{\rm comp}(\mathbf{x},0),$
for all $\mathbf{x}\in\Omega$.
Remark 8.
While our numerical experiments indicate that the iterative sequence $\{\mathbf{U}^{(k)}\}_{k\geq 0}$ converges reliably to the target solution when initialized from zero, we do not provide a theoretical justification of this convergence. The main difficulty lies in the lack of an appropriate Carleman estimate tailored to the anisotropic Navier–Stokes equations in the time-reduced framework. Developing such an estimate is highly nontrivial due to the complex structure of the fourth-order viscosity tensor. In the isotropic setting, where the viscosity tensor assumes a simpler form, one could potentially employ Carleman weight functions to construct and analyze the convergence of an analogous iterative scheme using existing Carleman estimates [9, 11]. However, since the focus of this work is on anisotropic media, we defer the theoretical convergence analysis for the isotropic case to future research.
5 Numerical study
In this section, we describe a semi-implicit finite difference scheme to generate simulated data, and the Picard quasi-reversibility method employed to solve the time-reduced model (3.17), thereby addressing Problem 1. We also present several numerical examples to illustrate the performance of the proposed approach.
5.1 Data generation
We present numerical simulations in the two-dimensional setting ($d=2$) for simplicity. The fourth-order viscosity tensor $\bm{\mu}$ used in our simulations is specified via its flattened representation:
$\bm{\mu}_{\rm flat}=\begin{bmatrix}\mu_{1111}&\mu_{1112}&\mu_{1121}&\mu_{1122}\\ \mu_{1211}&\mu_{1212}&\mu_{1221}&\mu_{1222}\\ \mu_{2111}&\mu_{2112}&\mu_{2121}&\mu_{2122}\\ \mu_{2211}&\mu_{2212}&\mu_{2221}&\mu_{2222}\end{bmatrix}=\begin{bmatrix}1&\tfrac{1}{16}&\tfrac{1}{16}&\tfrac{3}{8}\\ \tfrac{1}{16}&\tfrac{1}{4}&\tfrac{1}{4}&0\\ \tfrac{1}{16}&\tfrac{1}{4}&\tfrac{1}{4}&0\\ \tfrac{3}{8}&0&0&\tfrac{1}{2}\end{bmatrix}$.
This tensor does not correspond to the isotropic case. In isotropic linear viscoelasticity, the fourth-order viscosity tensor has the structure
$\mu_{ijkl}^{\rm iso}=\alpha\,\delta_{ij}\delta_{kl}+\theta\,(\delta_{ik}\delta_{jl}+\delta_{il}\delta_{jk}),$
where $\alpha$ and $\theta$ are scalar constants. In such a case, certain components, such as $\mu_{1211}^{\rm iso}$, are identically zero. However, for the tensor defined above, we have $\mu_{1211}=\tfrac{1}{16}\neq 0$, thereby violating the isotropic form. This confirms that the chosen $\bm{\mu}$ represents an anisotropic medium. For simplicity, in this numerical section, we choose $\nabla p=0$ as the pressure gradient does not influence the formulation of our reconstruction method.
We employ a semi-implicit finite difference scheme to solve the forward problem numerically. The computational domain $\Omega=(-1,1)^{2}$ is discretized using a uniform $41\times 41$ Cartesian grid, corresponding to $\Delta x=\Delta y=0.05$. The temporal interval $[0,T]$ with $T=1$ is uniformly partitioned into 700 time steps $\{t^{1},t^{2},\dots,t^{700}\}$, with the step size $\Delta t\approx 0.00143$.
For the time discretization, we denote $\mathbf{u}(x,y,t^{n+1})$ by ${\bf u}^{n}$ and adopt a semi-implicit scheme in which the time derivative is approximated by $\frac{\mathbf{u}^{n+1}-\mathbf{u}^{n}}{\Delta t}.$ The diffusion term involving the fourth-order viscosity operator ${\rm div}(\bm{\mu}:\nabla\mathbf{u})$ is treated implicitly at time $t^{n+1}$, while the nonlinear convection term $(\mathbf{u}\cdot\nabla)\mathbf{u}$ is computed explicitly using the known velocity field $\mathbf{u}^{n}$ from the previous time step. This leads to a linear system for ${\bf u}^{n+1}$ at each time step, which is solved using a sparse direct solver. Homogeneous Dirichlet boundary conditions are imposed at each step. This implicit-explicit like approach improves stability while keeping the computational complexity manageable.
After obtaining the numerical solution $\mathbf{u}^{*}$ to the forward problem using the semi-implicit finite difference method described above, we extract the associated Neumann boundary data $\partial_{\nu}\mathbf{u}^{*}$ by computing the normal derivative of the velocity field along the boundary $\partial\Omega$. This exact boundary flux is then used to generate synthetic measurement data by introducing multiplicative noise. Specifically, we define the noisy observations $\mathbf{f}=\mathbf{f}^{*}(1+\delta\,\mathcal{R})$, where $\mathbf{f}^{*}=\partial_{\nu}\mathbf{u}^{*}$, $\delta=10\%$ is the noise level, and $\mathcal{R}$ is a uniformly distributed random field with values in [1,1][-1,1]. This formulation reflects realistic measurement errors and serves to validate the robustness of the proposed inversion algorithm under noisy boundary conditions.
5.2 Numerical examples
With the boundary data available, we proceed to solve the inverse problem. The key computational steps of our proposed method, based on the time dimension reduction framework, are outlined in Algorithm 1.
In Step 1 of Algorithm 1, we determine the set of parameters through a trial-and-error process, guided by the results of a reference test case. Specifically, we choose $N=30$ and $\epsilon=10^{-6}$. After fine-tuning these parameters to achieve optimal performance in the reference test, we use the same set of parameters for all subsequent tests to ensure consistency and comparability across different scenarios.
In Step 4, we compute ${\bf U}^{(k+1)}$ based on the current knowledge of ${\bf U}^{(k)}$. As stated earlier, ${\bf U}^{(k+1)}$ is the solution to the linear system in (4.1), which can be obtained through least-squares optimization. The minimization process, once discretized, can be performed using optimization packages available in Matlab. In our implementation, we use the lsqlin function, a built-in tool in Matlab, to solve the problem. For further details on the implementation of this Carleman-weighted least-squares optimization with lsqlin, we refer to [21, 26].
Other steps of Algorithm 1 can be implemented straightforwardly.
Test 1.
In the first test, the true initial velocity field $\mathbf{U}_{0}^{\rm true}=\begin{bmatrix}u_{0}^{{\rm true},1}&u_{0}^{{\rm true},2}\end{bmatrix}^{\top}$ is constructed with high contrast between the regions of non-zero and background zero. Specifically, we set
$u_{0}^{{\rm true},1}(x,y)=\begin{cases}4,&\text{if }8(x-0.4)^{2}+y^{2}<0.64,\\ 0,&\text{otherwise},\end{cases}\quad u_{0}^{{\rm true},2}(x,y)=\begin{cases}4,&\text{if }x^{2}+8(y-0.3)^{2}<0.64,\\ 0,&\text{otherwise}.\end{cases}$
These profiles represent two elliptical regions: one vertically elongated in $u_{0}^{{\rm true},1}$ and the other horizontally elongated in $u_{0}^{{\rm true},2}$. This configuration allows us to test our method’s ability to recover both the magnitude and geometry of anisotropic features in the initial data, with high contrast between the active and inactive regions. The data is evolved forward under the anisotropic Navier–Stokes equations, and the resulting boundary observations are then used as input for the inverse reconstruction algorithm.
(a) $u_{0}^{{\rm true},1}$
(b) $u_{0}^{{\rm true},2}$
(c) $\ln\Big(\frac{\|{\bf U}^{(k)}-{\bf U}^{(k-1)}\|_{L^{\infty}}}{\|{\bf U}^{(k)}\|_{L^{\infty}}}\Big)$
(d) $u_{0}^{{\rm comp},1}$
(e) $u_{0}^{{\rm comp},2}$
Figure 1: Test 1: The true and computed solutions to the inverse initial data problem. The first row shows the true velocity fields and the logarithm of the error between consecutive iterations. The second row displays the computed velocity fields corresponding to the reconstruction.
The results presented in Figure 1 indicate a promising outcome for solving the inverse initial data problem, even in the presence of 10% noise in the data. The true velocity fields, shown in the first row, exhibit the expected localized features with distinct directional characteristics. The reconstructed velocity fields in the second row closely match the true data, demonstrating the effectiveness of the proposed method in accurately recovering the key structural elements of the original velocity field. From a quantitative perspective, the computed maximum value of $u_{0}^{{\rm comp},1}$ within the vertical ellipse is 4.1074 (with a relative error of 2.69%), and the computed maximum value of $u_{0}^{{\rm comp},2}$ within the horizontal ellipse is 4.0594 (with a relative error of 1.49%). Furthermore, the convergence plot (figure c) shows a consistent reduction in the error over successive iterations, emphasizing the method’s robustness. These results suggest that the method can provide reliable reconstructions, even with noisy observations, and holds potential for further refinement in more complex applications. Overall, the findings offer a positive outlook for the applicability of this approach in addressing real-world inverse problems effectively.
Test 2
For the second test case, the true initial velocity field consists of two components with localized features in the shape of diagonal ellipses. Specifically, the first component, $u_{0}^{\rm true,1}$, is defined as:
$u_{0}^{\rm true,1}(x,y)=\begin{cases}1,&\text{if }1.5(x+y)^{2}+y^{2}<0.7^{2},\\ 0,&\text{otherwise}.\end{cases}$
The second component, $u_{0}^{\rm true,2}$, takes the form:
$u_{0}^{\rm true,2}(x,y)=\begin{cases}1,&\text{if }x^{2}+1.5(x-y)^{2}<0.7^{2},\\ 0,&\text{otherwise}.\end{cases}$
Both components feature diagonal ellipses with distinct orientations, which serve as a test for the method’s ability to recover such geometrically oriented structures.
(a) $u_{0}^{{\rm true},1}$
(b) $u_{0}^{{\rm true},2}$
(c) $\ln\Big(\frac{\|{\bf U}^{(k)}-{\bf U}^{(k-1)}\|_{L^{\infty}}}{\|{\bf U}^{(k)}\|_{L^{\infty}}}\Big)$
(d) $u_{0}^{{\rm comp},1}$
(e) $u_{0}^{{\rm comp},2}$
Figure 2: Test 2: The true and computed solutions to the inverse initial data problem. The first row shows the true velocity fields and the logarithm of the error between consecutive iterations. The second row displays the computed velocity fields corresponding to the reconstruction.
The results displayed in Figure 2 demonstrate the effectiveness of the proposed method for solving the inverse initial data problem, even with 10% noise in the data. The true initial velocity fields, shown in the first row, exhibit diagonal elliptical structures with distinct orientations, accurately reflecting the expected features of the synthetic test case. In the second row, the computed velocity fields closely match the true data, highlighting the method’s ability to recover the key geometrical characteristics of the initial conditions despite the noisy observations. From a quantitative perspective, the computed maximum value of $u_{0}^{{\rm comp},1}$ within the ellipse is 0.8258 (with a relative error of 17.42%), and the computed maximum value of $u_{0}^{{\rm comp},2}$ within the ellipse is 0.8608 (with a relative error of 13.92%). These results show that the method is capable of capturing both the magnitude and shape of the velocity components, with errors remaining within an acceptable range. The logarithmic error plot in figure (c) illustrates the steady convergence of the method, providing further evidence of its robustness and accuracy in reconstructing the true data despite the presence of noise.
Test 3
In this test, the true initial velocity field is defined as follows. The first component, $u_{0}^{\rm true,1}(x,y)$, represents the characteristic function of an annular region. This is given by:
$u_{0}^{\rm true,1}(x,y)=\begin{cases}1,&\text{if }0.6^{2}<x^{2}+y^{2}<0.9^{2},\\ 0,&\text{otherwise}.\end{cases}$
The second component, $u_{0}^{\rm true,2}(x,y)$, is the characteristic function of a square with a central void, which is expressed as:
$u_{0}^{\rm true,2}(x,y)=\begin{cases}1,&\text{if }0.6<\max(|x|,|y|)<0.9,\\ 0,&\text{otherwise}.\end{cases}$
(a) $u_{0}^{{\rm true},1}$
(b) $u_{0}^{{\rm true},2}$
(c) $\ln\Big(\frac{\|{\bf U}^{(k)}-{\bf U}^{(k-1)}\|_{L^{\infty}}}{\|{\bf U}^{(k)}\|_{L^{\infty}}}\Big)$
(d) $u_{0}^{{\rm comp},1}$
(e) $u_{0}^{{\rm comp},2}$
Figure 3: Test 3: The true and computed solutions to the inverse initial data problem. The first row shows the true velocity fields and the logarithm of the error between consecutive iterations. The second row displays the computed velocity fields corresponding to the reconstruction.
The results displayed in Figure 3 highlight the effectiveness of our reconstruction method, even in the presence of 10% noise in the data. Despite the complexity of the true initial velocity field, which includes structures such as a ring-like formation and a square with a void, the reconstructed velocity fields (shown in the second row) closely match the true data (first row). This demonstrates the method’s ability to capture key structural features of the initial condition. The convergence plot in Figure (c) shows that the error between consecutive iterations consistently decreases, further emphasizing the robustness of the method. Quantitatively, the computed maximum value of $u_{0}^{{\rm comp},1}$ within the ring is 0.8756 (with a relative error of 12.44%), and the computed maximum value of $u_{0}^{{\rm comp},2}$ within the voided square is 0.9876 (with a relative error of 1.24%). These findings strongly suggest that the method holds significant promise for solving inverse initial data problems in real-world applications, even when the data is noisy.
6 On the stability of the time reduction and the choice of time basis
We detail how temporal projection and truncation act as a low-pass filter, justify the choice of the exponentially weighted Legendre basis, and discuss a scaled weight for large $T$.
6.1 Why time-dimensional reduction mitigates ill-posedness
It is well known that the inverse problem under consideration is ill-posed: small perturbations in the data can induce arbitrarily large perturbations in the reconstruction. We now explain how the time-dimensional reduction helps to reduce the degree of ill-posedness in practice, as demonstrated by the acceptable numerical examples above.
Recall that in the time-dimensional reduction framework, we only need the “low oscillation” time-projected boundary data
$\mathbf{f}_{m}(\mathbf{x})\;=\;\partial_{\nu}\mathbf{u}_{m}(\mathbf{x})\;=\;\int_{0}^{T}e^{-2t}\,\mathbf{f}(\mathbf{x},t)\,\Psi_{m}(t)\,dt,\qquad m=0,1,\dots,N,$
while the higher oscillation components when $m>N$ do not contribute to the computational procedure. Hence, when the measured data $\mathbf{f}$ contains noise, this projection-truncation step acts as a temporal low-pass filter. In other words, we only feed into the inversion those time components of the data that survive the weighted projection onto the first $N$ modes. This averaging and truncation reduce the effective noise level in the data-misfit and, hence, relax the practical severity of the ill-posedness.
6.2 Why use the exponentially weighted time basis
In this paper, we set $\Psi_{n}(t)=e^{t}Q_{n}(t)$, where $\{Q_{n}\}_{n\geq 0}$ are (shifted) Legendre polynomials. Although the factor $e^{t}$ “cancels” in the $L^{2}_{e^{-2t}}(0,T)$ inner product, it is essential when time derivatives appear. Indeed,
$\Psi_{n}^{\prime}(t)=\frac{d}{dt}\big(e^{t}Q_{n}(t)\big)=e^{t}\big(Q_{n}(t)+Q_{n}^{\prime}(t)\big),$
which is not identically zero on $(0,T)$. Consequently, in expressions such as (3.14), every coefficient ${\bf u}_{n}$ contributes through the term $\sum_{n=0}^{N}{\bf u}_{n}(\mathbf{x})\,\Psi_{n}^{\prime}(t)$, and this information propagates into the time-reduced model (3.15).
By contrast, using the unweighted Legendre basis $\{Q_{n}\}_{n\geq 0}$ gives $Q_{0}^{\prime}(t)=0$ for all $t$, as $Q_{0}$ is a constant. The corresponding coefficient ${\bf u}_{0}$ can be underrepresented in the derivative equations, which may weaken coupling between modes and degrade numerical conditioning. The exponential weighting avoids this loss by ensuring that even the lowest time mode remains visible to the derivative terms. This observation is confirmed by recomputing Test 1 with the time-dimensional reduction method using the unweighted basis $\{Q_{n}\}_{n\geq 0}$ in $L^{2}(0,T)$ in place of $\{\Psi_{n}\}_{n\geq 0}$; as shown in Figure 4, the exponentially weighted basis $\{\Psi_{n}\}_{n\geq 0}$ yields visibly better reconstructions.
(a) $u_{0}^{{\rm true},1}$
(b) $u_{0}^{{\rm comp},1}$ with the non weighted basis $\{Q_{n}\}_{n\geq 0}$
(c) $u_{0}^{{\rm comp},1}$ with the weighted basis $\{\Psi_{n}\}_{n\geq 0}$, see Figure 1d
(d) $u_{0}^{{\rm true},2}$
(e) $u_{0}^{{\rm comp},2}$ with the non weighted basis $\{Q_{n}\}_{n\geq 0}$
(f) $u_{0}^{{\rm comp},2}$ with the weighted basis $\{\Psi_{n}\}_{n\geq 0}$, see Figure 1e
Figure 4: Test 1: effect of the time basis in the time-dimensional reduction. (a,d) Ground-truth initial fields $u_{0}^{\mathrm{true},1}$ and $u_{0}^{\mathrm{true},2}$. (b,e) Reconstructions using the unweighted Legendre basis $\{Q_{n}\}_{n\geq 0}$ in $L^{2}(0,T)$. (c,f) Reconstructions using the exponentially weighted basis $\{\Psi_{n}\}_{n\geq 0}$ with $\Psi_{n}=e^{t}Q_{n}$ in $L^{2}_{e^{-2t}}(0,T)$. Under identical NN and regularization, the weighted basis yields sharper localization and amplitudes with fewer artifacts than the unweighted alternative.
Figure 4 compares reconstructions in Test 1 using the unweighted Legendre basis $\{Q_{n}\}_{n\geq 0}$ versus the exponentially weighted basis $\{\Psi_{n}\}_{n\geq 0}$ with $\Psi_{n}=e^{t}Q_{n}$ in $L^{2}_{e^{-2t}}(0,T)$. Figures (b,e) show that the unweighted choice produces blurrier shapes and lower peak amplitudes, with noticeable smearing away from the target. In contrast, figures (c,f) obtained with the weighted basis recover crisper boundaries, better peak levels, and fewer artifacts under the same $N$ and regularization. The improvement is consistent with the role of the factor $e^{t}$ in derivative terms: $\Psi_{n}^{\prime}(t)=e^{t}(Q_{n}(t)+Q_{n}^{\prime}(t))$ keeps the lowest time mode visible to $\partial_{t}$ and strengthens the coupling among modes in the time-dimensional reduction. Empirically, this yields a better conditioned inversion and more faithful reconstructions, which supports our choice of the weight function $e^{t}$.
6.3 On the exponential weight for large $T$
A natural concern is that the factor $e^{t}$ can become very large when $T$ is large. To avoid this, we may replace $e^{t}$ by a scaled weight $e^{\alpha t}$ with $\alpha=1/T$ (or more generally $\alpha=c/T$ for a modest constant $c$) and work in $L^{2}_{e^{-2\alpha t}}(0,T)$. In this case $\max_{t\in[0,T]}e^{\alpha t}=e^{\alpha T}=e^{c}$, so the weight remains uniformly bounded as $T$ increases. The construction of the time-dimensional reduction and all subsequent steps in this paper remain the same under this scaling; only the value of $\alpha$ is adjusted to control the dynamic range.
7 Concluding Remarks
This work presents a new computational approach to the inverse initial data problem for the anisotropic Navier–Stokes equations by leveraging a Legendre-based time reduction technique. By projecting the velocity field onto an exponentially weighted Legendre basis, we reformulated the original time-dependent problem into a sequence of time-independent elliptic systems, making the problem more tractable both analytically and numerically. This dimensional reduction not only simplifies the temporal complexity of the inverse problem but also enables stable recovery of initial velocity fields from partial and noisy boundary measurements.
The numerical results confirm that the proposed method yields accurate and stable reconstructions across diverse test cases, including those with high-contrast and geometrically complex initial conditions. Importantly, the technique exhibits robustness against significant levels of noise, demonstrating its practical potential for real-world applications where direct access to initial states is infeasible.
An important and challenging open question is whether the sequence $\{{\bf U}^{(k)}\}_{k\geq 0}$, defined in Section 4, converges to the true solution of the time-reduced model (3.17). A positive resolution of this question would mark a significant contribution to the theory of inverse problems. One powerful approach to establishing convergence involves incorporating Carleman weight functions into the cost functional $J^{\epsilon,N}_{{\bf U}^{(k)}}$ and utilizing Carleman estimates. However, appropriate Carleman estimates tailored for the anisotropic setting are not currently available.
If we instead restrict our attention to the isotropic case, specifically, the Lamé system, then existing Carleman estimates become applicable, see e.g., [9, 11]. In this context, some recently developed techniques, called the Carleman convexification [12, 14, 15, 19], the Carlaman contraction principle [20, 21, 27] and the Carleman–Newton method [1, 22] can be extended to construct a globally convergent sequence $\{{\bf U}^{(k)}\}_{k\geq 0}$ that approximates the true solution to the time-reduction model (3.17). In this case, both analytical and numerical global solvability of Problem 1 are achievable. Extending these ideas to the anisotropic setting forms a central direction for our future research.